Maher Corporation, which has only one product, has provided the following data concerning its most recent month of operations:

Selling price $ 198
Units in beginning inventory 0
Units produced 3,230
Units sold 3,060
Units in ending inventory 170
Variable costs per unit:
Direct materials $ 55
Direct labor $ 55
Variable manufacturing overhead $ 14
Variable selling and administrative expense $ 13
Fixed costs:
Fixed manufacturing overhead $ 109,820
Fixed selling and administrative expense $ 12,240

Required:

a. What is the unit product cost for the month under variable costing?

b. What is the unit product cost for the month under absorption costing?

c. Prepare a contribution format income statement for the month using variable costing.

d. Prepare an income statement for the month using absorption costing.

e. Reconcile the variable costing and absorption costing net operating incomes for the month.
Answer #1

Answer:

(a/b). Calculation of unit product cost for the month under variable costing and absorption costing method:

Particulars Variable costing method Absorption costing method
Direct Material $55 $55
Direct Labor $55 $55
Variable Manufacturing Overhead $14 $14
Fixed manufacturing overhead - $34*
Variable cost per unit $114 $148

*Fixed manufacturing cost per unit will be considered for absorption costing method: 109,820/3,230= $34

c. Calculation of income for the month under variable costing method:

Particulars Amount (in $)
Sales (3060*198) 605,880
Less:
Variable manufacturing costs (3060*114) 348,840
Variable selling costs (3060*13) 39,780
Contribution Margin: 217,260
Less:
Fixed manufacturing costs 109,820
Fixed Selling costs 12,240
Net Income for the month 95,200

d.  Calculation of income for the month under absorption costing method:

Particulars Amount (in $)
Sales (3060*198) 605,880
Less:
Cost of goods sold (3060*148) 452,880
Selling and administration costs (3060*16.78)* 51,376
Net Income for the month 101,624

*Selling cost per unit are calculated as follows: 13 per unit given + 12,240/3230: $16.78

e. Income under absorption costing method: $101,624

Income under variable costing method : $95,200

Reconciliation of above: Fixed cost included in income statement of Variable costing: 109,820+12,240=122,060

Fixed cost included in income statement of absorption costing costing: (109,820+12,240)*3,060/3230 = $115,636

(Please note that some figures are rounded off in solution)